An apparatus is designed to study insects at an acceleration of magnitude 980 m/sec2

(  100g ) The apparatus consists of a 2.0 m long rod with insect containers at either end.

The rod rotates about an axis perpendicular to the rod and at its center. Assume that the insects move at a constant speed (undergo uniform circular motion).

(a) How fast does an insect move when it experiences a radial acceleration of 980 m/sec2?

(b) What is the angular speed of the insect?
(c) How many revolutions do they undergo in one minute?
Answer #1

A)

Radius, r = 2/2 = 1 m

Acceleration, a = v^2/r

Velocity, v = sqrt(ar) = sqrt(980 x 1) = 31.31 m/s

B)

Angular velocity, w = v/r = 31.31/1 = 31.31 rad/sec

C)

Revolution = 31.31 x 60/2 pi = 298.94 rpm

Comment in case any doubt please rate my answer....